OpenStudy (anonymous):

Combine as indicated by the signs.4+a/a-3+4/5 i need help

OpenStudy (anonymous):

lcd is 5(a-3) = 20 + 5a + 4a - 12 ------------------- 5(a-3) = 9a + 8 -------- 5(a-3)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sure thing(:

OpenStudy (anonymous):

thanks